OpenStudy (anonymous):

An inclined plane is 5.00 m long and 3.00 m high. What is the ideal mechanical advantage of this machine? 1.00 , 2.00, 1.67 , or 2.50 ? :/

OpenStudy (michele_laino):

here is my reasoning:  in order to lift the box, if I use the inclined plane, I apply the subsequent force: \[F = mg\sin \theta \]

OpenStudy (anonymous):

:) yes

OpenStudy (michele_laino):

whereas if I don't use the inclibed plane I apply this force: \[R = mg\] namely the weight of the box 

OpenStudy (michele_laino):

so the advantage is given by the ratio R/F, namely: \[\frac{R}{F} = \frac{{mg}}{{mg\sin \theta }} = \frac{1}{{\sin \theta }}\]

OpenStudy (anonymous):

ok:)

OpenStudy (michele_laino):

now, what is sin(\theta) ?

OpenStudy (michele_laino):

it is simple, here is its value: \[H = L\sin \theta ,\quad \sin \theta = \frac{H}{L}\]

OpenStudy (michele_laino):

where H is the height of the inclined plane, where L is its length

OpenStudy (michele_laino):

so, replacing that expression into the formula for R/F, we get: \[\frac{R}{F} = \frac{{mg}}{{mg\sin \theta }} = \frac{1}{{\sin \theta }} = \frac{L}{H}\]

OpenStudy (anonymous):

okay! so what do we plug in? :/

OpenStudy (yadu123):

lenght and height

OpenStudy (michele_laino):

here is the next step: \[\frac{R}{F} = \frac{L}{H} = \frac{5}{3}\]

OpenStudy (anonymous):

ohh so we get 1.66666 so our solution is 1.67?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay! thank you!

OpenStudy (michele_laino):

:)